### v1

# How does Alignment Enhance LLMs’ Multilingual Capabilities? A Language Neurons Perspective

Abstract
Multilingual Alignment is an effective and representative paradigm to enhance LLMs’ multilingual capabilities, which transfers the capabilities from the high-resource languages to the low-resource languages. Meanwhile, some researches on language-specific neurons reveal that there are language-specific neurons that are selectively activated in LLMs when processing different languages. This provides a new perspective to analyze and understand LLMs’ mechanisms more specifically in multilingual scenarios. In this work, we propose a new finer-grained neuron identification algorithm, which detects language neurons (including language-specific neurons and language-related neurons) and language-agnostic neurons. Furthermore, based on the distributional characteristics of different types of neurons, we divide the LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting. Additionally, we systematically analyze the models before and after alignment with a focus on different types of neurons. We also analyze the phenomenon of “Spontaneous Multilingual Alignment”. Overall, our work conducts a comprehensive investigation based on different types of neurons, providing empirical results and valuable insights for better understanding multilingual alignment and multilingual capabilities of LLMs. [^1] The project will be available at https://github.com/NJUNLP/Language-Neurons-Alignment.
1 Introduction
By training on the extensive corpus, large language models (LLMs) demonstrate outstanding language capabilities (Grattafiori et al., 2024; Yang et al., 2024; Liu et al., 2024; Zhang et al., 2025). However, due to the unbalanced pretraining corpus across different languages, LLMs have very uneven performance on high-resource languages and low-resource languages (Huang et al., 2023; Zhu et al., 2023; Zhang et al., 2024). Therefore, researchers have conducted comprehensive explorations to further enhance the multilingual performance of LLMs. A straightforward approach is increasing the proportion of non-English texts during pretraining (Ni et al., 2021; Yang et al., 2024) or performing continual pretraining with multilingual texts (Liu et al., 2021; Ji et al., 2024). But these approaches often entail high computational costs and substantial amounts of multilingual data.
Considering LLMs’ great performance on the high-resource languages, multilingual alignment has emerged as a representative paradigm for enhancing multilingual capabilities by transferring knowledge from high-resource to low-resource languages (Zhao et al., 2024a; She et al., 2024). A representative example is MAPO She et al. (2024), which improves multilingual alignment by utilizing a well-trained multilingual translation model to compute alignment scores based on the conditional generation probability of translating non-English responses into English.
Figure 1: A neuron’s activation probability across different languages. This neuron that exhibits high activation probabilities across multiple but not all languages can’t be correctly categorized under the existing methodology.
Many studies conduct systematic mechanism analyses of the multilingual alignment and LLMs’ multilingual capabilities. Zhao et al. (2024b) split the multilingual processing workflow into three parts: multilingual understanding, resolving tasks, and generating outputs in the target language. This three-stage inference workflow clearly demonstrates how LLMs leverage English as a pivot language to handle multilingualism using a unified pattern. Inspired by the neurobiological underpinnings of human language faculties, Tang et al. (2024) conducts a fine-grained identification by detecting the language-specific neurons. Their results indicate that LLMs predominantly utilize a small subset of neurons to process the particular target language. Furthermore, these language-specific neurons are primarily situated in the model’s top and bottom layers (Tang et al., 2024), which is consistent with the three-stage multilingual workflow of Zhao et al. (2024b). However, we notice a key limitation in the existing language-specific neuron identification methodology: Some neurons are shared across multiple languages but are not entirely language-agnostic. Such neurons are incorrectly categorized as either language-specific or language-agnostic neurons under the existing framework. We present a case study in Figure 1. Furthermore, we aim to systematically explore an important question: Can we better analyze and understand how multilingual alignment enhances the LLMs’ multilingual capabilities from the perspective of language neurons?
In this work, we comprehensively investigate the multilingual alignment of LLMs from the perspective of language neurons, where we deploy MAPO as a representative multilingual alignment algorithm. Considering the above limitation of the existing language-specific neuron identification methodology, we define language neurons as the union of language-specific neurons and language-related neurons, as opposed to language-agnostic neurons. We separately distinguish language-related neurons from both language-specific and language-agnostic neurons, which allows for more precise analyses. Furthermore, we propose a new language neuron identification algorithm in our work, which is able to identify language-related neurons that are in fact shared across multiple languages.
Then we analyze the models before and after alignment, focusing on the changes in different types of neurons. Based on the distributional characteristics of these neurons, we divide LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting. Our findings reveal that different parts exhibit distinct dependencies on different types of neurons, and that multilingual alignment significantly enhances the activation of the corresponding types of neurons across the relevant layers. Additionally, we analyze the important “Spontaneous Multilingual Alignment” (Zhang et al., 2024) phenomenon in LLMs, providing insights into the roles of language-agnostic neurons and language-related neurons shared across languages. For further analysis, we also provide observations about the uniqueness of English and the neuron distributions. Overall, based on different types of neurons, we present empirical results and valuable insights that contribute to a deeper understanding of multilingual alignment and the multilingual capabilities of LLMs.
2 Related Work
2.1 Multilingual Alignment
Conducting pretraining or continual pretraining on the multilingual corpus is a straightforward and effective method to enhance LLMs’ multilingual capabilities (Ni et al., 2021; Ji et al., 2024). However, these methods typically require substantial investments in time, data, and computational resources. Thus, many researchers perform multilingual alignment to improve LLMs’ multilingual performance by transferring the capabilities from high-resource languages to low-resource languages (Eronen et al., 2023; Zhao et al., 2024c, a; She et al., 2024), which efficiently and effectively improves the model performance in low-resource language scenarios. Furthermore, Zhang et al. (2024) first finds the “Spontaneous Multilingual Alignment” phenomenon in LLMs, which demonstrates that conducting multilingual alignment based on a small number of languages effectively improves the alignment even between English and many languages unseen during alignment.
2.2 Mechanistic Interpretability
In addition to enhancing LLMs’ multilingual performance, research on the underlying mechanisms of multilingual capabilities in LLMs is still ongoing. It is crucial for us to understand and explain the LLMs and related methods explicitly. Typically, the existing approaches primarily perform mechanistic interpretability analyses by observing the internal states of the model (Nostalgebraist, 2020; Zhang et al., 2024; Zhao et al., 2024b; Mousi et al., 2024). Overall, neuron states and latent intermediate logits are both important objects of observation. For latent logits, Wendler et al. (2024) utilizes logit lens (Nostalgebraist, 2020) to directly project the logits in the intermediate layers to the vocabulary space, which reveals the latent participation of English in the intermediate layers. For neuron states, Hu et al. (2024) analyzes the neuron activation overlap to measure the extent of shared neuron activation across different languages.
2.3 Language-Specific Neurons
Many studies have revealed the language-related and language-agnostic components in LLMs. At the layer level, the multilingual processing of LLMs is considered to involve three stages (Zhao et al., 2024b; Wendler et al., 2024): converting multilingual inputs into a shared semantic space, intermediate-layer reasoning, and outputting in the target language. The top and bottom layers of the model handle multilingual processing, while the intermediate layers perform inference in similar patterns across different languages. This demonstrates a distinct division of labor within the model at the layer level regarding language specificity.
Furthermore, many studies investigate the finer-grained methods for language-specific neuron identification (Kojima et al., 2024; Tang et al., 2024). Tang et al. (2024) categorizes activated neurons into language-specific neurons and language-agnostic neurons. They detect language-specific neurons by calculating language activation probability entropy on massive text. However, we find that some neurons are activated by multiple languages (i.e., not language-specific), yet are not universally activated across all languages (i.e., not language-agnostic). Simply categorizing activated neurons into two classes blurs this distinction. Thus, we propose a new method to identify neurons, which categorizes activated neurons into three types: language-specific, language-related, and language-agnostic.
3 Methodology
In this section, we introduce our overall analysis pipeline in our work. First, we review the existing multilingual alignment algorithm and neuron analysis techniques as a preliminary study in §3.1. Then we introduce the multilingual alignment algorithm we utilize in our work in §3.2. Finally, for mechanistic interpretability analysis, we introduce our new method for detecting language-specific, language-related, and language-agnostic neurons (§3.3).
3.1 Preliminary Study
Most LLMs are pretrained mainly on the high-resource language corpus, which leads to LLMs’ unstable and unbalanced performance in multilingual scenarios. As a representative multilingual alignment algorithm, Multilingual-Alignment-as-Preference Optimization (MAPO) (She et al., 2024) effectively and efficiently improves the LLMs’ multilingual performance. Additionally, it is also important for us to understand and analyze the mechanism of LLMs’ multilingual capabilities and multilingual alignment. Moreover, some studies on the identification of the language-specific and language-agnostic neurons in LLMs (Tang et al., 2024; Kojima et al., 2024). It is found that LLMs’ capabilities of processing a particular language mainly come from a small subset of neurons (Tang et al., 2024).
However, there are still many important questions waiting for further investigation. On the one hand, many methods overlook neurons activated by multiple languages yet not language-agnostic, namely language-related neurons that lie between language-specific and language-agnostic categories. On the other hand, research from the perspective of language neurons on the underlying mechanisms of LLMs’ multilingual alignment and multilingual capabilities remains quite limited, which is essential for better understanding and improving the multilingual performance of LLMs.
3.2 Multilingual Alignment
MAPO is a typical multilingual alignment algorithm to align the reasoning capabilities of non-English language responses with those of English, which serves as the pivot language. Specifically, for a given query $X$ in a target (non-English) language and its corresponding English variant $X_{Eng}$, we collect their respective responses $Y$ and $Y_{Eng}$. An off-the-shelf translation model, parameterized by $\theta$, is deployed to estimate the conditional generation probability $P(Y\mid Y_{Eng};\theta)$ by force-decoding $Y$ conditioned on $Y_{Eng}$. A higher conditional probability is interpreted as stronger alignment between the target language response and its English counterpart. This probability is then used as an alignment score, denoted $r_{\theta}(X,Y)$.
This alignment score can be integrated into preference optimization algorithms. For instance, in PPO (Schulman et al., 2017), $r_{\theta}(X,Y)$ can be directly employed as the reward score. In DPO (Rafailov et al., 2023), for each target language, $n$ distinct outputs are generated. Based on the alignment score  these $n$ outputs are used to form $\binom{n}{2}$ preference pairs $(Y_{w},Y_{l})$, where $Y_{w}$ is deemed superior to $Y_{l}$ due to a higher alignment score. The model is then optimized by Eq1.
$L_{DPO}(\pi_{\theta};\pi_{ref})=-E_{(X,Y_{w},Y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(Y_{w}|X)}{\pi_{ref}(Y_{w}|X)}-\beta\log\frac{\pi_{\theta}(Y_{l}|X)}{\pi_{ref}(Y_{l}|X)}\right)\right]$ (1)
3.3 Language Neurons Identification
Following Tang et al. (2024), the neurons in our work are defined as a linear transformation of a single column in a weighted matrix $W$ followed by a non-linear activation, SiLU (Shazeer, 2020). For the $j$-th neuron in the $i$-th layer, its activation probability when processing responses in language $k$ is computed as:
$p_{i,j}^{k}=\mathbb{E}\left(\mathbb{I}\left(\mathrm{SiLU}(x^{i}W^{i})_{j}>0\right)\mid\text{language}\ k\right)$ (2)
We define language neurons as those exhibiting higher activation probabilities for some languages, while discriminating from others. To select neurons that, despite exhibiting relatively higher entropy, still demonstrate high activation probabilities for some languages, we additionally incorporate the maximum activation probability, as formulated in Eq 3:
$\mathrm{score}_{i,j}=-\sum_{k=1}^{l}p_{i,j}^{\prime k}\,\log p_{i,j}^{\prime k}\;-\;\lambda\,\max_{1\leq k\leq l}p_{i,j}^{k},$ (3)
where $p_{i,j}^{\prime}$ represents the probability distribution $p_{i,j}$ after normalization and $\lambda$ is a balancing coefficient. Neurons with scores falling in the lowest percentile, specifically, the bottom 5% are selected.
Furthermore, to identify how many languages each selected neuron is related to, we introduce a threshold $\tau$ and compute:
$N_{i,j}=\sum_{k=1}^{l}\mathbb{I}\bigl{(}p_{i,j}^{k}>\tau\bigr{)}.$ (4)
A neuron is considered as a language-specific neuron if $N_{i,j}=1$, and as a language-related neuron if $1<N_{i,j}<l$. Both of them belong to language neuron. In contrast, a neuron is considered language-agnostic neuron if it exhibits high activation probabilities across all $l$ languages.
Finally, given our focus on multilingual reasoning tasks, we select neurons exclusively based on responses from multilingual reasoning datasets, rather than relying on multilingual plain text (Tang et al., 2024).
4 Experiments
4.1 Experimental Setup
Following She et al. (2024), we conduct our experiments and analyses on the mathematical reasoning tasks and different languages. In this section, we introduce our experimental settings in detail.
Models
We include two different models in our experiments and analyses. Following She et al. (2024), we conduct our experiments on MistralMathOctopus-7B  [^2] https://huggingface.co/kevinpro/MistralMathOctopus-7B and MetaMathOctopus-7B  [^3] https://huggingface.co/kevinpro/MetaMathOctopus-7B. MistralMathOctopus is obtained by fine-tuning MetaMath-Mistral (Yu et al., 2023) with MGSM8KInstruct (Chen et al., 2023). MetaMathOctopus is obtained by fine-tuning MetaMath (Yu et al., 2023) with MGSM8KInstruct. Considering limited computational resources and reproducibility, we directly utilize the publicly released base models. Our analyses are mainly based on MistralMathOctopus in the main text and we report more results in the Appendix.
Datasets
We conduct experiments on two representative mathematical reasoning benchmarks, MGSM (Shi et al., 2022) and MSVAMP (Chen et al., 2023). MGSM is a widely used benchmark for multilingual mathematical reasoning evaluation. MSVAMP is an out-of-domain test set in contrast to MGSM, which evaluates robustness and generalization (Zhu et al., 2024; She et al., 2024).
Languages
Following She et al. (2024), we choose the following 10 different languages for analysis in our work. As a pivot language, English (en) is used as the alignment target. We also choose Chinese (zh), Russian (ru), German (de), French (fr), Spanish (es), Japanese (ja), Swahili (sw), Thai (th) and Bengali (bn) as 9 representative non-English languages.
(a) Base - Language-Specific Neurons
(b) Aligned - Language-Specific Neurons
(c) Base - Language Neurons
(d) Aligned - Language Neurons
Figure 2: PPL changes of MistralMathOctopus on MGSM after deactivating language-specific neurons or language neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model. For comparison, the results of Tang et al. (2024) are provided in Appendix D.
Implementations
Due to limited computational resources, our exploration focuses on the most effective DPO variant of MAPO (She et al., 2024). We select 1, 4, and 8 tasks from the NumGLUE (Mishra et al., 2022), an arithmetic reasoning benchmark, and translate questions into 9 languages, consistent with the MGSM, thereby creating a multilingual seed dataset. To construct preference pairs, we sample responses using the corresponding base models and employ NLLB-200-distilled-600M [^4] https://huggingface.co/facebook/nllb-200-distilled-600M as the translation model to obtain alignment scores. Finally, for each model and each target language (excluding English), we gain 10,000 preference pairs. Training is conducted using LoRA (Hu et al., 2022). During the neuron selection stage, we perform force-decoding on the responses of the MGSM or MSVAMP dataset to obtain the activation probabilities of neurons for each language. Based on empirical results on development sets, we set the balancing coefficient $\lambda=0.04$ and the threshold $\tau=0.5$. Additional implementation details are provided in Appendix B.
4.2 Language Neurons Identification
Based on the neuron identification algorithm introduced in §3.3, we identify the language-specific neurons, language-related neurons, and language-agnostic neurons in the model. To further validate the effectiveness of our algorithm, we follow  Tang et al. (2024) by examining changes in the perplexity of LLMs by deactivating the identified language neurons across different languages. Experiments are conducted on both the base model and the aligned model, with results presented in Figure 2. We report the results of both language-specific neurons and language neurons.
It can be found that whether deactivating language-specific neurons or all language neurons, the results consistently exhibit the same pattern: the diagonal elements in each row show the highest values. Notably, deactivating language neurons leads to a more pronounced effect compared to deactivating only language-specific neurons. These observations support the following findings: (1) Our algorithm effectively identifies language-specific and language-related neurons; (2) For a given language, in addition to its language-specific neurons, there are also a substantial number of shared language-related neurons contributing to its performance; (3) Deactivating all the language-related neurons of one language doesn’t cause significant impacts on the model’s performance in other languages. The above findings confirm the validity of the language neurons identified by our method and further provide insights into the characteristics of language neurons.
Figure 3: Layer-wise distribution of the different types of neurons of MistralMathOctopus on MGSM.
4.3 Layer-wise Functionality Analysis
Based on the identified neurons, we perform layer-wise functional analyses of all layers in the LLMs. We begin by analyzing the distributions of different types of neurons in the base model. And we report the results in Figure 3. Through the analysis of the distribution of different types of neurons, we can further divide the LLMs’ internal process for multilingual inference into four parts rather than the three-stage division discussed in some works (Wendler et al., 2024; Zhao et al., 2024b):
  1. 1.
    Multilingual Understanding: In the initial layers, the number of language neurons (language-specific and language-related neurons) peaks, while the number of language-agnostic neurons is relatively low. The model maps multilingual inputs into a unified semantic space at this stage.
  2. 2.
    Shared Semantic Space Reasoning: In the intermediate layers, the model engages in reasoning within a shared semantic space across different languages. During this stage, language neurons are largely absent, whereas language-agnostic neurons become dominant.
  3. 3.
    Multilingual Output Space Transformation: The model transfers features into the multilingual output space in this stage in preparation for generating the final output. In this part, the number of language neurons reaches a peak again, while the number of language-agnostic neurons drops to the lowest point.
  4. 4.
    Vocabulary Space Outputting: In the last layer, the model maps vectors of different languages into a shared vocabulary space to generate outputs. The number of both language-related and language-agnostic neurons rises sharply, while the language-specific neurons are fewer than those in several previous layers.
Meanwhile, the distribution of different types of neurons aligns with the conclusions from existing studies mentioned above. Overall, we can find that the number of neurons varies correspondingly with the different inference stages of LLMs.
4.4 Layer-wise Neuron Changes Analysis
In §4.3, we investigate the fundamental distribution of neurons and the basic partitioning of the model. We further analyze the changes in different types of neurons before and after multilingual alignment. Based on the four functional stages in LLMs, we quantify the layer-wise changes ($\Delta$) in the number of different types of neurons. Figure 4 presents the results for language-specific neurons, language-related neurons, and language-agnostic neurons.
Figure 4: Layer-wise changes in the number of different types of neurons of MistralMathOctopus on MGSM.
During the multilingual understanding stage, the number of language neurons increases, while language-agnostic neurons decrease. In the subsequent shared semantic space reasoning stage, language-agnostic neurons increase substantially, whereas language neurons remain stable and nearly absent. Then, in the third stage, as language-agnostic neurons decrease, language neurons increase overall. Additionally, we notice that the number of language-related neurons show an upward trend. Finally, in the last stage, the number of language-agnostic increases significantly in the aligned model, accompanied by a reduction in language neurons. We also report the results of different checkpoints during the alignment process in Appendix F.
Overall, we find that language neurons and language-agnostic neurons exhibit generally opposite trends across different layers, which corresponds to the characteristics of LLMs at different stages of inference. Especially, at the last stage, language-agnostic neurons play an more important role than language neurons. Multilingual alignment facilitates more effective activation of the appropriate neurons at each stage, thereby improving the model’s capability to handle multilingual tasks.
4.5 Macroscopic Analysis of Different Types of Neurons
We further conduct macroscopic analysis for different types of neurons. In our neuron identification algorithm introduced in §3.3, the number of languages that share a specific neuron is an attribution characterizing all types of activated neurons. Since our study involves 10 languages, the valid range of $N$ is from $1$ to $10$. Among these, values of $N$ from $2$ to $9$ correspond to language-related neurons. As special cases in our work, $N=1$ represents language-specific neurons, while $N=10$ corresponds to language-agnostic neurons.
Figure 5: Changes in the number of neurons shared by $N$ languages after alignment.
We report the changes in the number of neurons after multilingual alignment for each value of $N$ ranging from 1 to 10 in Figure 5. The results show a decrease in the number of language-specific neurons, while an increase in the number of language-related neurons, which are shared across multiple languages. This indicates that multilingual alignment encourages LLMs to develop and utilize more shared language-related neurons, rather than on language-specific neurons, which are applicable to only a single language. Meanwhile, during the alignment process, the model improves its understanding of task-relevant common knowledge. Therefore, the overall number of language-agnostic neurons also increases significantly.
Table 1: Accuracy of the MistralMathOctopus base model and aligned model on MGSM. "X/Y $\Rightarrow$ T" indicates that languages X and Y are used for multilingual alignment.
MGSM bn th sw ja zh ru de es fr en Avg.
base 43.6 53.2 50.4 55.6 59.6 59.2 61.2 62.8 56.8 75.6 57.8
zh/de $\Rightarrow$ en 46.4 55.6 59.2 56.8 64.0 71.2 66.8 71.2 69.2 75.2 63.6
sw/th $\Rightarrow$ en 48.8 58.8 59.2 56.4 68.4 68.4 69.2 69.6 70.4 77.6 64.7
4.6 Spontaneous Multilingual Alignment Analysis
The important “Spontaneous Multilingual Alignment” phenomenon is first revealed and discussed by Zhang et al. (2024), which means that conducting alignment in a small number of languages significantly improves multilingual alignment even between English and many languages unseen during the alignment process. We further analyze this phenomenon in our experiments. As shown in Table 1, spontaneous multilingual alignment also emerges under the multilingual alignment strategy employed in our study. Except for the languages used for alignment, LLMs exhibit notable performance gain in other unaligned languages.
To understand how multilingual alignment generalizes to other languages, we analyze the changes in different types of neurons before and after multilingual alignment based on our method.
Table 2: Average results of neuron count changes across multiple languages. “Trained” indicates the trained languages in the spontaneous multilingual alignment experiment. “Others” indicates other languages except the trained languages. We round the results to the nearest integer.
Language Language-Specific Language-Related
Trained -37 +232
Others -36 +205
Taking the case of “zh/de $\Rightarrow$ en” as a representative example, we report the average results in Table 4.6. For the trained languages, the number of language-specific neurons decreases, while the number of language-related neurons increases. This indicates that the aligned languages tend to utilize more language-related neurons shared with other languages rather than exclusive language-specific neurons. Moreover, we extend this analysis to languages other than the trained languages and observe a similar phenomenon. These findings indicate that multilingual alignment facilitates the use of language-related neurons while reducing the reliance on language-specific neurons in both trained and other unseen languages. We hypothesize that the new language-related neurons shared with trained languages contribute to the performance improvement on other unseen languages.
Table 3: Average number of different types of neurons for English and non-English languages of MistralMathOctopus on MGSM. We round the results to the nearest integer.
Language Language-Specific Language-Related
English 46 603
non-English 613 2006
Table 4: Overlap ratio of different types of neurons across different domains and before and after alignment. Following She et al. (2024), MSVAMP is regarded as an out-of-domain dataset. The results of MistralMathOctopus on MGSM are used as the fiducial value.
Variable (%) Language-Specific Language-Related
Domain 80.7 92.3
Alignment 95.6 92.1
4.7 Further Analysis
Uniqueness of English
Since current LLMs are primarily pretrained on English data, English is often regarded as playing a special role within LLMs (Wendler et al., 2024). In our experiments, we also observe that English exhibits markedly different characteristics compared to other non-English languages. Based on the identified neurons in our work, in Figure 2, we can find that deactivating the language neurons of English has a negligible impact on the model’s performance in English, which is entirely different from the behavior observed in other languages. This is also consistent with the results of Tang et al. (2024). Furthermore, based on this finding, we quantify the number of language neurons for English and non-English languages based on the MistralMathOctopus base model (Table 4.6). Our analysis reveals that English has significantly fewer neurons than other languages, both in terms of language-specific and language-related neurons. We hypothesize that this is due to the fact that English actually possesses numerous language-related neurons. And since English serves as a pivot language, these language-related neurons are likely shared with almost all other languages. It causes them to be confounded with language-agnostic neurons.
Stability of Neuron Distributions
We discuss the stability of neuron distributions across different data domains, as well as before and after alignment. To quantify the stability of neuron distributions, we compute the neuron overlap ratio in both settings, with the results summarized in Table 4.6. We can find that although the exact positions of a few language neurons may vary across different settings, the positional distribution of most language neurons remains stable. This also indicates good reliability and generalization of the language neurons identified under fixed hyperparameters.
5 Conclusion
In this work, we systematically investigate the multilingual alignment from the perspective of language neurons. We propose a new language neuron identification algorithm based on entropy and probability value, which detects the language-specific neurons, language-related neurons, and language-agnostic neurons in LLMs. The validity of the identified neurons is confirmed through deactivation ablation experiments. Furthermore, we examine the multilingual alignment mechanism by analyzing the roles of different types of neurons. Based on their distributional characteristics, we categorize LLMs’ internal process into four functional parts. Our analysis reveals that multilingual alignment enhances the model’s utilization of the corresponding types of neurons across different functional parts. Meanwhile, we find that alignment promotes a greater reliance on shared language-related neurons across languages, rather than on language-specific neurons. We also explore the phenomenon of spontaneous multilingual alignment. Additionally, we provide further analysis and more empirical results based on the preceding findings.
Appendix A Limitations
We provide insights and analysis results for multilingual alignment and multilingual capabilities of LLMs, which allows for a better understanding of multilingualism. Despite we have conducted a systematic investigation in our work, there are still some limitations waiting for research. Due to the limited resources, we only conduct experiments on two different models and two different datasets. We are willing to perform a more comprehensive analysis on different scenarios if more resources are available in the future. Additionally, we don’t perform a finer-grained analysis of neurons within the same layer in this work. We would like to explore this in our future work.
Appendix B Implementation Details
Our experiments are conducted on 4 NVIDIA RTX A6000 GPUs or 4 NVIDIA GeForce RTX 3090 GPUs. We use the TRL [von Werra et al., 2020] and DeepSpeed [Rasley et al., 2020] frameworks for preference alignment, and the vLLM engine Kwon et al. [2023] for inference.
B.1 MAPO
We employ the officially released scripts [^5] https://github.com/NJUNLP/MAPO to generate the preference data. For the alignment process, we set the learning rate to 1e-6 and the batch size to 16. LoRA is utilized to fine-tune the model with a LoRA rank of 64, a LoRA alpha of 128, and a LoRA dropout rate of 0.05. The total number of training steps is set to 1000. It takes 7.5 hours for one alignment.
B.2 Language Neurons Identification
We spend 25 minutes calculating the activation probability of each neuron with MGSM dataset on a single NVIDIA GeForce RTX 3090 GPU. We spend 2 hours calculating the activation probability of each neuron with MSVAMP dataset on a single NVIDIA GeForce RTX 3090 GPU. After obtaining the activation probability, it takes 3 minutes to identify how many languages each selected neuron is related to.
Appendix C Licenses for Used Assets
We list the names of the licenses for each asset we utilized in our work:
  • MGSM: CC-BY-SA-4.0
  • MSVAMP: Apache-2.0
  • GSM8KInstruct: Apache-2.0
  • NumGLUE: Apache-2.0
  • MistralMathOctopus: Apache-2.0
  • MetaMathOctopus: Apache-2.0
  • NLLB-200-distilled-600M: CC-BY-NC-4.0
  • TRL: Apache-2.0
  • DeepSpeed: Apache-2.0
  • vLLM: Apache-2.0
Appendix D Deactivation Ablation Experiments
For the MistralMathOctopus model, Figure 6 presents the results of the deactivation ablation experiments conducted on MGSM using the neuron identification algorithm proposed by Tang et al. [2024]. Additionally, Figure 7 presents the results of the deactivation ablation experiments on the out-of-domain test set, MSVAMP. To verify the generalization of our identification algorithm, we also deploy the MetaMathOctopus to conduct the deactivation ablation experiment, with the result in Figure 8.
Appendix E Layer-wise distribution of the different types of neurons
In Figure 9(a) and 9(b), we separately display the layer-wise distribution of the different types of neurons of MistarlMathOctopus on MSVAMP and MetaMathOctopus on MGSM to validate the generalization of our finding discussed in §4.3.
Appendix F Layer-wise Changes in the Number of Different Types of Neurons During Alignment
[…]
Appendix G Spontaneous Multilingual Alignment
We report the accuracy of the base model on the MGSM and MSVAMP benchmarks in Table 5. The expected accuracy of random guessing is 50.0%. "X/Y $\Rightarrow$ T" indicates that languages X and Y are used for alignment. The best results for each language are highlighted. It is evident that models trained on multilingual translation data substantially outperform the original model across a wide range of languages, indicating that multilingual training significantly enhances the model’s multilingual capabilities.
(a) Base - Language-Specific Neurons
(b) Aligned - Language-Specific Neurons
Figure 6: PPL changes of MistralMathOctopus on MGSM when deploying the algorithm proposed by Tang et al. [2024].
(a) Base - Language-Specific Neurons
(b) Aligned - Language-Specific Neurons
(c) Base - Language Neurons
(d) Aligned - Language Neurons
Figure 7: PPL changes of MistralMathOctopus on MSVAMP after deactivating language-specific neurons or language neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model.
(a) Base - Language Neurons on MGSM
(b) Aligned - Language Neurons on MGSM
Figure 8: PPL changes of MetaMathOctopus on MGSM after deactivating language neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model.
(a) MistralMathOctopus on MSVAMP.
(b) MetaMathOctopus on MGSM.
Figure 9: Layer-wise distribution of the different types of neurons.
Figure 10: Layer-wise changes in the number of MetaMathOctopus
(a) MGSM
(b) MSVAMP
Figure 11: Layer-wise changes in the number of language-specific neurons of MistralMathOctopus during the alignment.
(a) MGSM
(b) MSVAMP
Figure 12: Layer-wise changes in the number of language-related neurons of MistralMathOctopus during the alignment.
(a) MGSM
(b) MSVAMP
Figure 13: Layer-wise changes in the number of language-agnostic neurons of MistralMathOctopus during the alignment.
Tested on MGSM bn th sw ja zh ru de es fr en Avg.
base 43.6 53.2 50.4 55.6 59.6 59.2 61.2 62.8 56.8 75.6 57.8
zh $\Rightarrow$ en 49.6 58.4 54.8 56.4 65.2 70.0 66.4 72.8 68.8 78.4 64.1
zh/de $\Rightarrow$ en 46.4 55.6 59.2 56.8 64.0 71.2 66.8 71.2 69.2 75.2 63.6
sw/th $\Rightarrow$ en 48.8 58.8 59.2 56.4 68.4 68.4 69.2 69.6 70.4 77.6 64.7
zh/es/ru $\Rightarrow$ en 46.0 56.4 58.8 54.8 63.2 70.8 68.8 71.6 69.6 76.8 63.7
zh/es/fr/ja/de/sw/ru/th/bn $\Rightarrow$ en 49.6 60.0 56.4 56.4 64.4 64.8 65.2 65.2 61.2 76.4 62.0
Tested on MSVAMP bn th sw ja zh ru de es fr en Avg.
base 49.3 62.5 60.6 60.9 67.4 64.9 66.5 67.6 67.2 77.0 64.4
zh $\Rightarrow$ en 52.8 62.5 60.9 63.7 66.6 67.5 69.4 69.8 69.7 76.5 65.9
zh/de $\Rightarrow$ en 53.2 62.0 62.9 65.3 66.1 68.7 69.9 69.0 70.1 77.0 66.4
sw/th $\Rightarrow$ en 53.8 67.4 65.1 67.7 71.9 71.4 71.9 72.0 72.8 78.9 69.3
zh/es/ru $\Rightarrow$ en 52.9 63.2 61.5 63.0 67.8 68.1 69.8 68.7 70.6 78.3 66.4
zh/es/fr/ja/de/sw/ru/th/bn $\Rightarrow$ en 54.9 65.9 65.6 68.6 69.7 70.2 72.3 71.6 71.6 77.3 68.8
Table 5: Accuracy of the base model and aligned variants on benchmarks.

### v2

@@
 Multilingual Alignment is an effective and representative paradigm to enhance LLMs’ multilingual capabilities, which transfers the capabilities from the high-resource languages to the low-resource languages.
-Meanwhile, some researches on language-specific neurons reveal that there are language-specific neurons that are selectively activated in LLMs when processing different languages.
-This provides a new perspective to analyze and understand LLMs’ mechanisms more specifically in multilingual scenarios.
-In this work, we propose a new finer-grained neuron identification algorithm, which detects language neurons (including language-specific neurons and language-related neurons) and language-agnostic neurons.
+Meanwhile, some research on language-specific neurons provides a new perspective to analyze and understand LLMs’ mechanisms.
+However, we find that there are many neurons that are shared by multiple but not all languages and cannot be correctly classified.
+In this work, we propose a ternary classification methodology that categorizes neurons into three types, including language-specific neurons, language-related neurons, and general neurons.
+And we propose a corresponding identification algorithm to distinguish these different types of neurons.
 Furthermore, based on the distributional characteristics of different types of neurons, we divide the LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting.
@@
 We also analyze the phenomenon of “Spontaneous Multilingual Alignment”.
-Overall, our work conducts a comprehensive investigation based on different types of neurons, providing empirical results and valuable insights for better understanding multilingual alignment and multilingual capabilities of LLMs.
-[^1] The project will be available at https://github.com/NJUNLP/Language-Neurons-Alignment.
-1 Introduction
-By training on the extensive corpus, large language models (LLMs) demonstrate outstanding language capabilities (Grattafiori et al., 2024; Yang et al., 2024; Liu et al., 2024; Zhang et al., 2025).
-However, due to the unbalanced pretraining corpus across different languages, LLMs have very uneven performance on high-resource languages and low-resource languages (Huang et al., 2023; Zhu et al., 2023; Zhang et al., 2024).
+Overall, our work conducts a comprehensive investigation based on different types of neurons, providing empirical results and valuable insights to better understand multilingual alignment and multilingual capabilities of LLMs.
+Codehttps://github.com/NJUNLP/Language-Neurons-Alignment
+Extended versionhttps://arxiv.org/abs/2505.21505
+Introduction
+By training on the extensive corpus, large language models (LLMs) demonstrate outstanding language capabilities (Yang et al.
+2024; Liu et al.
+2024; Grattafiori et al.
+2024; Zhang et al.
+2025).
+However, due to the unbalanced pretraining corpus across different languages, LLMs have very uneven performance on high-resource languages and low-resource languages (Huang et al.
+2023; Zhu et al.
+2023; Zhang et al.
+2024; Fan et al.
+2025).
 Therefore, researchers have conducted comprehensive explorations to further enhance the multilingual performance of LLMs.
-A straightforward approach is increasing the proportion of non-English texts during pretraining (Ni et al., 2021; Yang et al., 2024) or performing continual pretraining with multilingual texts (Liu et al., 2021; Ji et al., 2024).
+A straightforward approach is increasing the proportion of non-English texts during pretraining (Ni et al.
+2021; Yang et al.
+2024) or performing continual pretraining with multilingual texts (Liu, Winata, and Fung 2021; Ji et al.
+2024).
 But these approaches often entail high computational costs and substantial amounts of multilingual data.
-Considering LLMs’ great performance on the high-resource languages, multilingual alignment has emerged as a representative paradigm for enhancing multilingual capabilities by transferring knowledge from high-resource to low-resource languages (Zhao et al., 2024a; She et al., 2024).
-A representative example is MAPO She et al.
-(2024), which improves multilingual alignment by utilizing a well-trained multilingual translation model to compute alignment scores based on the conditional generation probability of translating non-English responses into English.
 Figure 1: A neuron’s activation probability across different languages.
-This neuron that exhibits high activation probabilities across multiple but not all languages can’t be correctly categorized under the existing methodology.
+This neuron that exhibits high activation probabilities across multiple (green) but not all (blue) languages can’t be correctly categorized under the existing classification methodology.
+The dashed line denotes the threshold that determines whether a neuron exhibits high activation in a given language.
+Considering LLMs’ great performance on high-resource languages, multilingual alignment has emerged as a representative paradigm for enhancing multilingual reasoning by transferring knowledge from high-resource to low-resource languages (Zhao et al.
+2024a; She et al.
+2024).
+A representative example is MAPO (Multilingual-Alignment-as-Preference Optimization) (She et al.
+2024), which improves multilingual alignment by utilizing a well-trained multilingual translation model to compute alignment scores based on the conditional generation probability of translating non-English responses into English.
 Many studies conduct systematic mechanism analyses of the multilingual alignment and LLMs’ multilingual capabilities.
@@
 Inspired by the neurobiological underpinnings of human language faculties, Tang et al.
-(2024) conducts a fine-grained identification by detecting the language-specific neurons.
-Their results indicate that LLMs predominantly utilize a small subset of neurons to process the particular target language.
-Furthermore, these language-specific neurons are primarily situated in the model’s top and bottom layers (Tang et al., 2024), which is consistent with the three-stage multilingual workflow of Zhao et al.
+(2024) categorize neurons in LLMs into two primary types: language-specific neurons and general neurons.
+Notably, these language-specific neurons are primarily situated in the model’s top and bottom layers (Tang et al.
+2024), which is consistent with the three-stage multilingual workflow of Zhao et al.
 (2024b).
-However, we notice a key limitation in the existing language-specific neuron identification methodology: Some neurons are shared across multiple languages but are not entirely language-agnostic.
-Such neurons are incorrectly categorized as either language-specific or language-agnostic neurons under the existing framework.
-We present a case study in Figure 1.
-Furthermore, we aim to systematically explore an important question: Can we better analyze and understand how multilingual alignment enhances the LLMs’ multilingual capabilities from the perspective of language neurons?
-In this work, we comprehensively investigate the multilingual alignment of LLMs from the perspective of language neurons, where we deploy MAPO as a representative multilingual alignment algorithm.
-Considering the above limitation of the existing language-specific neuron identification methodology, we define language neurons as the union of language-specific neurons and language-related neurons, as opposed to language-agnostic neurons.
-We separately distinguish language-related neurons from both language-specific and language-agnostic neurons, which allows for more precise analyses.
-Furthermore, we propose a new language neuron identification algorithm in our work, which is able to identify language-related neurons that are in fact shared across multiple languages.
+However, we identify a key limitation in the existing neuron classification methodology: the identification of language-specific neurons focuses on the processing a particular language and thus neglects the inter-language alignment; while general neurons encode universal knowledge which is independent of a specific language.
+This leads to a critical question: what is the mechanism of sharing neurons between some languages?
+Furthermore, it is not well understood how multilingual alignment enhances the LLMs’ multilingual reasoning capabilities from the perspective of language neurons.
+In this work, we comprehensively investigate the multilingual alignment of LLMs with MAPO (She et al.
+2024) as a representative multilingual alignment algorithm.
+As shown in Figure 1, we observed the existence of neurons that exhibit high activation probability on multiple but not all languages.
+These neurons are neither language-specific nor general for all languages.
+Hence, we refer to them as language-related neurons.
+Moreover, we redefine language-specific neurons to restrict each neuron to be activated for only one language.
+The general neurons are defined as the neurons that are effective for all languages.
+To facilitate subsequent analysis, we propose a corresponding identification algorithm for distinguishing these different types of neurons.
 Then we analyze the models before and after alignment, focusing on the changes in different types of neurons.
-Based on the distributional characteristics of these neurons, we divide LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting.
-Our findings reveal that different parts exhibit distinct dependencies on different types of neurons, and that multilingual alignment significantly enhances the activation of the corresponding types of neurons across the relevant layers.
-Additionally, we analyze the important “Spontaneous Multilingual Alignment” (Zhang et al., 2024) phenomenon in LLMs, providing insights into the roles of language-agnostic neurons and language-related neurons shared across languages.
+Based on the distributional characteristics of these neurons, we divide LLMs’ internal process for multilingual inference into four parts, with different parts exhibit distinct dependencies on different types of neurons.
+We demonstrate that multilingual alignment significantly enhances the activation of the corresponding types of neurons across the relevant layers.
+Additionally, we analyze the “spontaneous multilingual alignment” (Zhang et al.
+2024) phenomenon in LLMs, providing insights into the roles of general neurons and language-related neurons shared across languages.
 For further analysis, we also provide observations about the uniqueness of English and the neuron distributions.
-Overall, based on different types of neurons, we present empirical results and valuable insights that contribute to a deeper understanding of multilingual alignment and the multilingual capabilities of LLMs.
-2 Related Work
-2.1 Multilingual Alignment
-Conducting pretraining or continual pretraining on the multilingual corpus is a straightforward and effective method to enhance LLMs’ multilingual capabilities (Ni et al., 2021; Ji et al., 2024).
+Overall, based on different types of neurons, we present empirical results and valuable insights that contribute to a deeper understanding of multilingual alignment and the multilingual reasoning capabilities of LLMs.
+Related Work
+Multilingual Alignment
+Conducting pretraining or continual pretraining on the multilingual corpus is a straightforward and effective method to enhance LLMs’ multilingual capabilities (Ni et al.
+2021; Ji et al.
+2024).
 However, these methods typically require substantial investments in time, data, and computational resources.
-Thus, many researchers perform multilingual alignment to improve LLMs’ multilingual performance by transferring the capabilities from high-resource languages to low-resource languages (Eronen et al., 2023; Zhao et al., 2024c, a; She et al., 2024), which efficiently and effectively improves the model performance in low-resource language scenarios.
+Thus, many researchers perform multilingual alignment to improve LLMs’ multilingual performance by transferring the capabilities from high-resource languages to low-resource languages (Eronen, Ptaszynski, and Masui 2023; Zhao et al.
+2024c, a; She et al.
+2024), which efficiently and effectively improves the model performance in low-resource language scenarios.
 Furthermore, Zhang et al.
 (2024) first finds the “Spontaneous Multilingual Alignment” phenomenon in LLMs, which demonstrates that conducting multilingual alignment based on a small number of languages effectively improves the alignment even between English and many languages unseen during alignment.
-2.2 Mechanistic Interpretability
+Mechanistic Interpretability
 In addition to enhancing LLMs’ multilingual performance, research on the underlying mechanisms of multilingual capabilities in LLMs is still ongoing.
 It is crucial for us to understand and explain the LLMs and related methods explicitly.
-Typically, the existing approaches primarily perform mechanistic interpretability analyses by observing the internal states of the model (Nostalgebraist, 2020; Zhang et al., 2024; Zhao et al., 2024b; Mousi et al., 2024).
+Typically, the existing approaches primarily perform mechanistic interpretability analyses by observing the internal states of the model (Nostalgebraist 2020; Zhang et al.
+2024; Zhao et al.
+2024b; Mousi et al.
+2024).
 Overall, neuron states and latent intermediate logits are both important objects of observation.
 For latent logits, Wendler et al.
-(2024) utilizes logit lens (Nostalgebraist, 2020) to directly project the logits in the intermediate layers to the vocabulary space, which reveals the latent participation of English in the intermediate layers.
+(2024) utilizes logit lens (Nostalgebraist 2020) to directly project the logits in the intermediate layers to the vocabulary space, which reveals the latent participation of English in the intermediate layers.
 For neuron states, Hu et al.
 (2024) analyzes the neuron activation overlap to measure the extent of shared neuron activation across different languages.
-2.3 Language-Specific Neurons
-Many studies have revealed the language-related and language-agnostic components in LLMs.
-At the layer level, the multilingual processing of LLMs is considered to involve three stages (Zhao et al., 2024b; Wendler et al., 2024): converting multilingual inputs into a shared semantic space, intermediate-layer reasoning, and outputting in the target language.
+Language-Specific Neurons
+Many studies have revealed the language-related and universal components in LLMs.
+At the layer level, the multilingual processing of LLMs is considered to involve three stages (Zhao et al.
+2024b; Wendler et al.
+2024): converting multilingual inputs into a shared semantic space, intermediate-layer reasoning, and outputting in the target language.
 The top and bottom layers of the model handle multilingual processing, while the intermediate layers perform inference in similar patterns across different languages.
 This demonstrates a distinct division of labor within the model at the layer level regarding language specificity.
-Furthermore, many studies investigate the finer-grained methods for language-specific neuron identification (Kojima et al., 2024; Tang et al., 2024).
+Furthermore, many studies investigate the finer-grained methods for language-specific neuron identification (Kojima et al.
+2024; Tang et al.
+2024; Tan, Wu, and Monz 2024).
 Tang et al.
-(2024) categorizes activated neurons into language-specific neurons and language-agnostic neurons.
+(2024) categorizes activated neurons into language-specific neurons and general neurons.
 They detect language-specific neurons by calculating language activation probability entropy on massive text.
-However, we find that some neurons are activated by multiple languages (i.e., not language-specific), yet are not universally activated across all languages (i.e., not language-agnostic).
+However, we find that some neurons are activated by multiple languages (i.e., not language-specific), yet are not universally activated across all languages (i.e., not general).
 Simply categorizing activated neurons into two classes blurs this distinction.
-Thus, we propose a new method to identify neurons, which categorizes activated neurons into three types: language-specific, language-related, and language-agnostic.
-3 Methodology
-In this section, we introduce our overall analysis pipeline in our work.
-First, we review the existing multilingual alignment algorithm and neuron analysis techniques as a preliminary study in §3.1.
-Then we introduce the multilingual alignment algorithm we utilize in our work in §3.2.
-Finally, for mechanistic interpretability analysis, we introduce our new method for detecting language-specific, language-related, and language-agnostic neurons (§3.3).
-3.1 Preliminary Study
+Thus, we propose a new method to identify neurons, which categorizes activated neurons into three types: language-specific neurons, language-related neurons, and general neurons.
+Methodology
+Preliminary Study
 Most LLMs are pretrained mainly on the high-resource language corpus, which leads to LLMs’ unstable and unbalanced performance in multilingual scenarios.
-As a representative multilingual alignment algorithm, Multilingual-Alignment-as-Preference Optimization (MAPO) (She et al., 2024) effectively and efficiently improves the LLMs’ multilingual performance.
+As a representative multilingual alignment algorithm, Multilingual-Alignment-as-Preference Optimization (MAPO) (She et al.
+2024) effectively and efficiently improves the LLMs’ multilingual performance.
 Additionally, it is also important for us to understand and analyze the mechanism of LLMs’ multilingual capabilities and multilingual alignment.
-Moreover, some studies on the identification of the language-specific and language-agnostic neurons in LLMs (Tang et al., 2024; Kojima et al., 2024).
-It is found that LLMs’ capabilities of processing a particular language mainly come from a small subset of neurons (Tang et al., 2024).
+Moreover, some studies on the identification of the language-specific and general neurons in LLMs (Tang et al.
+2024; Kojima et al.
+2024).
+It is found that LLMs’ capabilities of processing a particular language mainly come from a small subset of neurons (Tang et al.
+2024).
 However, there are still many important questions waiting for further investigation.
-On the one hand, many methods overlook neurons activated by multiple languages yet not language-agnostic, namely language-related neurons that lie between language-specific and language-agnostic categories.
-On the other hand, research from the perspective of language neurons on the underlying mechanisms of LLMs’ multilingual alignment and multilingual capabilities remains quite limited, which is essential for better understanding and improving the multilingual performance of LLMs.
-3.2 Multilingual Alignment
+On the one hand, many methods overlook neurons activated by multiple languages yet not general, namely language-related neurons lie between language-specific and general categories.
+On the other hand, research from the perspective of language neurons on the underlying mechanisms of LLMs’ multilingual alignment and multilingual reasoning capabilities remains quite limited, which is essential for better understanding and improving the multilingual performance of LLMs.
+Multilingual Alignment
 MAPO is a typical multilingual alignment algorithm to align the reasoning capabilities of non-English language responses with those of English, which serves as the pivot language.
-Specifically, for a given query $X$ in a target (non-English) language and its corresponding English variant $X_{Eng}$, we collect their respective responses $Y$ and $Y_{Eng}$.
-An off-the-shelf translation model, parameterized by $\theta$, is deployed to estimate the conditional generation probability $P(Y\mid Y_{Eng};\theta)$ by force-decoding $Y$ conditioned on $Y_{Eng}$.
+Specifically, for a given query $X$ in a target (non-English) language and its corresponding English variant $X_{En}$, we collect their respective responses $Y$ and $Y_{En}$.
+An off-the-shelf translation model, parameterized by $\theta$, is deployed to estimate the conditional generation probability $P(Y\mid Y_{En};\theta)$ by force-decoding $Y$ conditioned on $Y_{En}$.
 A higher conditional probability is interpreted as stronger alignment between the target language response and its English counterpart.
@@
 This alignment score can be integrated into preference optimization algorithms.
-For instance, in PPO (Schulman et al., 2017), $r_{\theta}(X,Y)$ can be directly employed as the reward score.
-In DPO (Rafailov et al., 2023), for each target language, $n$ distinct outputs are generated.
-Based on the alignment score  these $n$ outputs are used to form $\binom{n}{2}$ preference pairs $(Y_{w},Y_{l})$, where $Y_{w}$ is deemed superior to $Y_{l}$ due to a higher alignment score.
-The model is then optimized by Eq1.
-$L_{DPO}(\pi_{\theta};\pi_{ref})=-E_{(X,Y_{w},Y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(Y_{w}|X)}{\pi_{ref}(Y_{w}|X)}-\beta\log\frac{\pi_{\theta}(Y_{l}|X)}{\pi_{ref}(Y_{l}|X)}\right)\right]$ (1)
-3.3 Language Neurons Identification
+In PPO (Schulman et al.
+2017), $r_{\theta}(X,Y)$ can be directly employed as the reward score.
+In DPO (Rafailov et al.
+2023), for each target language, $n$ distinct outputs are generated.
+Based on the alignment score, these $n$ outputs are used to form $\binom{n}{2}$ preference pairs $(Y_{w},Y_{l})$, where $Y_{w}$ is deemed superior to $Y_{l}$ due to a higher alignment score.
+The model is optimized by Eq.
+1 to Eq.
+3:
+$A$ $=\beta\log\frac{\pi_{\theta}(Y_{w}|X)}{\pi_{ref}(Y_{w}|X)}$ (1)
+$B$ $=\beta\log\frac{\pi_{\theta}(Y_{l}|X)}{\pi_{ref}(Y_{l}|X)}$ (2)
+$L_{DPO}(\pi_{\theta};\pi_{ref})$ $=-E_{(X,Y_{w},Y_{l})\sim D}\left[\log\sigma\left(A-B\right)\right]$ (3)
+(a) Language-specific Neurons
+(b) Language-specific & Language-related Neurons
+Figure 2: PPL changes of MistralMathOctopus on MGSM after deactivating language-specific neurons or language-specific & language-related neurons.
+For comparison, the results of Tang et al.
+(2024) are provided in appendix.
+Language Neurons Identification
 Following Tang et al.
-(2024), the neurons in our work are defined as a linear transformation of a single column in a weighted matrix $W$ followed by a non-linear activation, SiLU (Shazeer, 2020).
+(2024), the neurons in our work are defined as a linear transformation of a single column in a weighted matrix $W$ followed by a non-linear activation, SiLU (Shazeer 2020).
 For the $j$-th neuron in the $i$-th layer, its activation probability when processing responses in language $k$ is computed as:
 $p_{i,j}^{k}=\mathbb{E}\left(\mathbb{I}\left(\mathrm{SiLU}(x^{i}W^{i})_{j}>0\right)\mid\text{language}\ k\right)$ (4)
-We define language neurons as those exhibiting higher activation probabilities for some languages, while discriminating from others.
-To select neurons that, despite exhibiting relatively higher entropy, still demonstrate high activation probabilities for some languages, we additionally incorporate the maximum activation probability, as formulated in Eq 3:
+We define language-related neurons as those neurons exhibiting high activation probabilities for multiple but not all languages.
+However, methods that only focus on language specificity are defective in detecting the above neurons.
+To better detect the different types of neurons, we trade off two intrinsic properties of each neuron: (1) Language-specificity, which is quantified by the entropy of its activation probability distribution across languages; (2) Effectiveness, which measures the extent to which the neuron participates when the model solves tasks.
+And it is quantified by the neuron’s maximum activation probability across different languages in our work.
+By doing so, we are better able to simultaneously identify neurons that are relatively more specific and relatively more generalized.
+They are combined into a unified metric as formulated in Eq 5:
 $\mathrm{score}_{i,j}=-\sum_{k=1}^{l}p_{i,j}^{\prime k}\,\log p_{i,j}^{\prime k}\;-\;\lambda\,\max_{1\leq k\leq l}p_{i,j}^{k},$ (5)
 where $p_{i,j}^{\prime}$ represents the probability distribution $p_{i,j}$ after normalization and $\lambda$ is a balancing coefficient.
-Neurons with scores falling in the lowest percentile, specifically, the bottom 5% are selected.
-Furthermore, to identify how many languages each selected neuron is related to, we introduce a threshold $\tau$ and compute:
+Specifically, we automatically determine a value of $\lambda$ such that general neurons are not identified as language-specific or language-related neurons shared by all $l$ languages.
+Following Tang et al.
+(2024), neurons with scores falling in the lowest $1\%$ are selected.
+Furthermore, to identify how many languages each selected neuron is related to, we introduce a threshold $\tau$:
 $N_{i,j}=\sum_{k=1}^{l}\mathbb{I}\bigl(p_{i,j}^{k}>\tau\bigr).$ (6)
+Following Tang et al.
+(2024), the threshold $\tau$ is set to the top 5% of all activation probabilities.
 A neuron is considered as a language-specific neuron if $N_{i,j}=1$, and as a language-related neuron if $1<N_{i,j}<l$.
-Both of them belong to language neuron.
-In contrast, a neuron is considered language-agnostic neuron if it exhibits high activation probabilities across all $l$ languages.
-Finally, given our focus on multilingual reasoning tasks, we select neurons exclusively based on responses from multilingual reasoning datasets, rather than relying on multilingual plain text (Tang et al., 2024).
-4 Experiments
-4.1 Experimental Setup
-Following She et al. (2024), we conduct our experiments and analyses on the mathematical reasoning tasks and different languages. In this section, we introduce our experimental settings in detail.
+Meanwhile, a neuron is considered general neuron if it exhibits high activation probabilities across all $l$ languages.
+Finally, given our focus on multilingual reasoning tasks, we select neurons exclusively based on responses from multilingual reasoning datasets, rather than relying on multilingual plain text (Tang et al.
+2024).
+Experiments
+Experimental Setup
+Following She et al. (2024), we conduct our experiments and analyzes on the mathematical reasoning tasks in different languages. In this section, we introduce our experimental settings in detail.
 Models
@@
 (2024), we conduct our experiments on MistralMathOctopus-7B  [^1] https://huggingface.co/kevinpro/MistralMathOctopus-7B and MetaMathOctopus-7B  [^2] https://huggingface.co/kevinpro/MetaMathOctopus-7B.
-MistralMathOctopus is obtained by fine-tuning MetaMath-Mistral (Yu et al., 2023) with MGSM8KInstruct (Chen et al., 2023).
-MetaMathOctopus is obtained by fine-tuning MetaMath (Yu et al., 2023) with MGSM8KInstruct.
+MistralMathOctopus is obtained by fine-tuning MetaMath-Mistral (Yu et al.
+2023) with MGSM8KInstruct (Chen et al.
+2023).
+MetaMathOctopus is obtained by fine-tuning MetaMath (Yu et al.
+2023) with MGSM8KInstruct.
 Considering limited computational resources and reproducibility, we directly utilize the publicly released base models.
@@
 Datasets
-We conduct experiments on two representative mathematical reasoning benchmarks, MGSM (Shi et al., 2022) and MSVAMP (Chen et al., 2023).
+We conduct experiments on two representative mathematical reasoning benchmarks, MGSM (Shi et al.
+2022) and MSVAMP (Chen et al.
+2023).
 MGSM is a widely used benchmark for multilingual mathematical reasoning evaluation.
-MSVAMP is an out-of-domain test set in contrast to MGSM, which evaluates robustness and generalization (Zhu et al., 2024; She et al., 2024).
+MSVAMP is an out-of-domain test set in contrast to MGSM, which evaluates robustness and generalization (Zhu et al.
+2024; She et al.
+2024).
 Languages
 Following She et al.
-(2024), we choose the following 10 different languages for analysis in our work.
+(2024), we choose the following 10 different languages for analysis.
 As a pivot language, English (en) is used as the alignment target.
 We also choose Chinese (zh), Russian (ru), German (de), French (fr), Spanish (es), Japanese (ja), Swahili (sw), Thai (th) and Bengali (bn) as 9 representative non-English languages.
-(a) Base - Language-Specific Neurons
-(b) Aligned - Language-Specific Neurons
-(c) Base - Language Neurons
-(d) Aligned - Language Neurons
-Figure 2: PPL changes of MistralMathOctopus on MGSM after deactivating language-specific neurons or language neurons.
-“Base” indicates the results of the base model.
-“Aligned” indicates the results of the aligned model.
-For comparison, the results of Tang et al.
-(2024) are provided in Appendix D.
+Figure 3: Layer-wise distribution of the different types of neurons of MistralMathOctopus on MGSM.
 Implementations
-Due to limited computational resources, our exploration focuses on the most effective DPO variant of MAPO (She et al., 2024).
-We select 1, 4, and 8 tasks from the NumGLUE (Mishra et al., 2022), an arithmetic reasoning benchmark, and translate questions into 9 languages, consistent with the MGSM, thereby creating a multilingual seed dataset.
+Due to limited computational resources, our exploration focuses on the most effective DPO variant of MAPO (She et al.
+2024).
+We select 1, 4, and 8 tasks from the NumGLUE (Mishra et al.
+2022), an arithmetic reasoning benchmark, and translate questions into 9 languages, consistent with the MGSM, thereby creating a multilingual seed dataset.
 To construct preference pairs, we sample responses using the corresponding base models and employ NLLB-200-distilled-600M [^3] https://huggingface.co/facebook/nllb-200-distilled-600M as the translation model to obtain alignment scores.
 Finally, for each model and each target language (excluding English), we gain 10,000 preference pairs.
-Training is conducted using LoRA (Hu et al., 2022).
+Training is conducted using LoRA (Hu et al.
+2022).
 During the neuron selection stage, we perform force-decoding on the responses of the MGSM or MSVAMP dataset to obtain the activation probabilities of neurons for each language.
-Based on empirical results on development sets, we set the balancing coefficient $\lambda=0.04$ and the threshold $\tau=0.5$.
-Additional implementation details are provided in Appendix B.
-4.2 Language Neurons Identification
-Based on the neuron identification algorithm introduced in §3.3, we identify the language-specific neurons, language-related neurons, and language-agnostic neurons in the model.
+Additional implementation details are provided in appendix.
+Language Neurons Identification
+Based on the neuron identification algorithm, we identify the language-specific neurons, language-related neurons, and general neurons in the model.
 To further validate the effectiveness of our algorithm, we follow  Tang et al.
-(2024) by examining changes in the perplexity of LLMs by deactivating the identified language neurons across different languages.
-Experiments are conducted on both the base model and the aligned model, with results presented in Figure 2.
-We report the results of both language-specific neurons and language neurons.
-It can be found that whether deactivating language-specific neurons or all language neurons, the results consistently exhibit the same pattern: the diagonal elements in each row show the highest values.
-Notably, deactivating language neurons leads to a more pronounced effect compared to deactivating only language-specific neurons.
-These observations support the following findings: (1) Our algorithm effectively identifies language-specific and language-related neurons; (2) For a given language, in addition to its language-specific neurons, there are also a substantial number of shared language-related neurons contributing to its performance; (3) Deactivating all the language-related neurons of one language doesn’t cause significant impacts on the model’s performance in other languages.
+(2024) by examining changes in the perplexity (PPL) scores of LLMs by deactivating the identified neurons across different languages.
+Experiments are conducted on the base model, with results presented in Figure 2.
+We report the results of two settings, deactivating language-specific neurons and deactivating language-specific & language-related neurons.
+The higher PPL change (darker cells) indicates the stronger reliance on the neurons be deactivated.
+It can be found that the performance for each language mostly relies on both its language-specific neurons and language-related neurons rather than other neurons, as the diagonal elements in each row show the highest changes in PPL.
+However, different languages vary in the extent to which they rely on their corresponding language-specific and language-related neurons, reflecting differences in cross-lingual alignment across different languages.
+Notably, deactivating both language-specific and language-related neurons leads to a more pronounced effect compared to deactivating only language-specific neurons.
+Suggesting that for a given language, in addition to its language-specific neurons, shared language-related neurons also significantly contribute to its performance.
+Also, deactivating all the language-related neurons of one language doesn’t cause significant impacts on the model’s performance in other languages, which indicates that the language-related neurons for a specific language are relatively evenly dispersed across multiple other languages.
 The above findings confirm the validity of the language neurons identified by our method and further provide insights into the characteristics of language neurons.
-Figure 3: Layer-wise distribution of the different types of neurons of MistralMathOctopus on MGSM.
-4.3 Layer-wise Functionality Analysis
+Layer-wise Functionality Analysis
 Based on the identified neurons, we perform layer-wise functional analyses of all layers in the LLMs.
@@
 And we report the results in Figure 3.
-Through the analysis of the distribution of different types of neurons, we can further divide the LLMs’ internal process for multilingual inference into four parts rather than the three-stage division discussed in some works (Wendler et al., 2024; Zhao et al., 2024b):
+Some works have divided the LLM’s multilingual inference process into three stages (Wendler et al.
+2024; Zhao et al.
+2024b).
+However, through the analysis of the distribution of different types of neurons, we find that although the combined function of the last stage is still to produce language-specific symbols, the behavior of language-specific neurons first reaches its peaks and then declines, which differs substantially from the other two types.
+So we further divide the final stage into two parts, leading to a four-stage interpretation of the LLMs’ internal process for multilingual inference:
   1. 1.
-    Multilingual Understanding: In the initial layers, the number of language neurons (language-specific and language-related neurons) peaks, while the number of language-agnostic neurons is relatively low.
+    Multilingual Understanding: In the initial layers, the number of both language-specific and language-related neurons peaks, while the number of general neurons is relatively low.
 The model maps multilingual inputs into a unified semantic space at this stage.
@@
     Shared Semantic Space Reasoning: In the intermediate layers, the model engages in reasoning within a shared semantic space across different languages.
-During this stage, language neurons are largely absent, whereas language-agnostic neurons become dominant.
+During this stage, both language-specific and language-related neurons are largely absent, whereas general neurons become dominant.
   3. 3.
     Multilingual Output Space Transformation: The model transfers features into the multilingual output space in this stage in preparation for generating the final output.
-In this part, the number of language neurons reaches a peak again, while the number of language-agnostic neurons drops to the lowest point.
+In this part, the number of both language-specific and language-related neurons reaches a peak again, while the number of general neurons drops to the lowest point.
   4. 4.
     Vocabulary Space Outputting: In the last layer, the model maps vectors of different languages into a shared vocabulary space to generate outputs.
-The number of both language-related and language-agnostic neurons rises sharply, while the language-specific neurons are fewer than those in several previous layers.
-Meanwhile, the distribution of different types of neurons aligns with the conclusions from existing studies mentioned above.
-Overall, we can find that the number of neurons varies correspondingly with the different inference stages of LLMs.
-4.4 Layer-wise Neuron Changes Analysis
-In §4.3, we investigate the fundamental distribution of neurons and the basic partitioning of the model.
-We further analyze the changes in different types of neurons before and after multilingual alignment.
-Based on the four functional stages in LLMs, we quantify the layer-wise changes ($\Delta$) in the number of different types of neurons.
-Figure 4 presents the results for language-specific neurons, language-related neurons, and language-agnostic neurons.
+The total number of language-specific neurons and language-related neurons increases, but there is an opposite trend between language-specific and language-related neurons.
+Additionally, unlike assumptions in previous work (Wendler et al.
+2024; Zhao et al.
+2024b), the number of general neurons reaches maxima at this stage, consistent with the behavior in the early layers.
+We therefore speculate that this may be related to the shared vocabulary across different languages.
+Meanwhile, the distribution of different types of neurons aligns with the conclusions from the existing studies mentioned above.
+Overall, the number of neurons varies correspondingly with the different inference stages of LLMs.
+Layer-wise Neuron Changes Analysis
+We further analyze the changes in different types of neurons before and after alignment.
+Based on the four functional stages above, we quantify the layer-wise changes ($\Delta$) in the number of different types of neurons in Figure 4.
 Figure 4: Layer-wise changes in the number of different types of neurons of MistralMathOctopus on MGSM.
-During the multilingual understanding stage, the number of language neurons increases, while language-agnostic neurons decrease.
-In the subsequent shared semantic space reasoning stage, language-agnostic neurons increase substantially, whereas language neurons remain stable and nearly absent.
-Then, in the third stage, as language-agnostic neurons decrease, language neurons increase overall.
-Additionally, we notice that the number of language-related neurons show an upward trend.
-Finally, in the last stage, the number of language-agnostic increases significantly in the aligned model, accompanied by a reduction in language neurons.
-We also report the results of different checkpoints during the alignment process in Appendix F.
-Overall, we find that language neurons and language-agnostic neurons exhibit generally opposite trends across different layers, which corresponds to the characteristics of LLMs at different stages of inference.
-Especially, at the last stage, language-agnostic neurons play an more important role than language neurons.
+During the multilingual understanding stage, the number of both language-specific and language-related neurons increases, while general neurons decrease.
+In the subsequent shared semantic space reasoning stage, general neurons increase substantially, whereas both language-specific neurons and language-related neurons remain stable and nearly absent.
+In the third stage, as general neurons decrease, the sum of language-specific neurons and language-related neurons increase overall.
+Additionally, we notice that language-specific and language-related neurons show opposite trends.
+This reflects that the aligned model relies more on shared language-related neurons than language-specific neurons in the third stage, which is difficult to observe under the previous taxonomies.
+Finally, in the last stage, the number of general neurons increases significantly in the aligned model, accompanied by a reduction in both language-specific and language-related neurons.
+This contradicts the conventional three-stage assumption, indicating that general neurons also play a significant role in the final vocabulary outputting stage.
+Additionally, combined with the observations in Figure 5, the model relies on a smaller set of language neurons with a higher degree of sharing in this stage.
+We also report the results of different checkpoints during the alignment process in appendix.
+Overall, we find the union set of language-specific and language-related neurons and general neurons exhibit generally opposite trends across layers, which corresponds to the characteristics of LLMs at different stages.
+Especially, at the last stage, general neurons play an more important role.
 Multilingual alignment facilitates more effective activation of the appropriate neurons at each stage, thereby improving the model’s capability to handle multilingual tasks.
-4.5 Macroscopic Analysis of Different Types of Neurons
+Macroscopic Analysis of Neurons
 We further conduct macroscopic analysis for different types of neurons.
-In our neuron identification algorithm introduced in §3.3, the number of languages that share a specific neuron is an attribution characterizing all types of activated neurons.
+In our neuron identification algorithm, the number of languages that share a specific neuron is an attribution characterizing all types of activated neurons.
 Since our study involves 10 languages, the valid range of $N$ is from $1$ to $10$.
 Among these, values of $N$ from $2$ to $9$ correspond to language-related neurons.
-As special cases in our work, $N=1$ represents language-specific neurons, while $N=10$ corresponds to language-agnostic neurons.
-Figure 5: Changes in the number of neurons shared by $N$ languages after alignment.
+As special cases in our work, $N=1$ represents language-specific neurons, while $N=10$ corresponds to general neurons.
+Figure 5: Changes in the number of neurons shared by NN languages after alignment.
 We report the changes in the number of neurons after multilingual alignment for each value of $N$ ranging from 1 to 10 in Figure 5.
 The results show a decrease in the number of language-specific neurons, while an increase in the number of language-related neurons, which are shared across multiple languages.
-This indicates that multilingual alignment encourages LLMs to develop and utilize more shared language-related neurons, rather than on language-specific neurons, which are applicable to only a single language.
-Meanwhile, during the alignment process, the model improves its understanding of task-relevant common knowledge.
-Therefore, the overall number of language-agnostic neurons also increases significantly.
-Table 1: Accuracy of the MistralMathOctopus base model and aligned model on MGSM. "X/Y $\Rightarrow$ T" indicates that languages X and Y are used for multilingual alignment.
+This indicates that multilingual alignment encourages LLMs to develop and utilize more shared language-related neurons, rather than language-specific neurons, which are applicable to only a single language.
+Meanwhile, during the alignment, the model improves its task-relevant common knowledge.
+Therefore, the number of general neurons also increases significantly.
+Language-related neurons and language-specific neurons exhibit overall opposite trends, further highlighting the importance of a more fine-grained analysis of the two categories within our new taxonomy.
+In addition, we also report the overall neuron numbers for each value of $N$ before and after alignment in Figure 10.
 MGSM bn th sw ja zh ru de es fr en Avg.
@@
 sw/th $\Rightarrow$ en 48.8 58.8 59.2 56.4 68.4 68.4 69.2 69.6 70.4 77.6 64.7
-4.6 Spontaneous Multilingual Alignment Analysis
-The important “Spontaneous Multilingual Alignment” phenomenon is first revealed and discussed by Zhang et al.
-(2024), which means that conducting alignment in a small number of languages significantly improves multilingual alignment even between English and many languages unseen during the alignment process.
+Table 1: Accuracy of the MistralMathOctopus base model and aligned model on MGSM. ”X/Y $\Rightarrow$ T” indicates that languages X and Y are used for multilingual alignment.
+Spontaneous Multilingual Alignment Analysis
+The “spontaneous multilingual alignment” phenomenon is first revealed and discussed by Zhang et al.
+(2024), which shows that conducting alignment in a small number of languages significantly improves multilingual alignment even between English and many languages unseen during the alignment process.
 We further analyze this phenomenon in our experiments.
-As shown in Table 1, spontaneous multilingual alignment also emerges under the multilingual alignment strategy employed in our study.
-Except for the languages used for alignment, LLMs exhibit notable performance gain in other unaligned languages.
+As shown in Table 1, spontaneous multilingual alignment also emerges under the multilingual alignment method employed in our study.
+Besides the languages used for alignment, LLMs exhibit notable performance gains in other unaligned languages.
 To understand how multilingual alignment generalizes to other languages, we analyze the changes in different types of neurons before and after multilingual alignment based on our method.
+Language Language-Specific Language-Related
+Trained -37 +232
+Others -36 +205
 Table 2: Average results of neuron count changes across multiple languages.
@@
 We round the results to the nearest integer.
-Language Language-Specific Language-Related
-Trained -37 +232
-Others -36 +205
-Taking the case of “zh/de $\Rightarrow$ en” as a representative example, we report the average results in Table 4.6.
+Taking the case of “zh/de $\Rightarrow$ en” as a representative example, we report the average results in Table 2.
 For the trained languages, the number of language-specific neurons decreases, while the number of language-related neurons increases.
-This indicates that the aligned languages tend to utilize more language-related neurons shared with other languages rather than exclusive language-specific neurons.
+This indicates the aligned languages tend to utilize more language-related neurons shared with other languages rather than exclusive language-specific neurons.
 Moreover, we extend this analysis to languages other than the trained languages and observe a similar phenomenon.
-These findings indicate that multilingual alignment facilitates the use of language-related neurons while reducing the reliance on language-specific neurons in both trained and other unseen languages.
+These findings indicate multilingual alignment facilitates the use of language-related neurons while reducing the reliance on language-specific neurons in both trained and other unseen languages.
 We hypothesize that the new language-related neurons shared with trained languages contribute to the performance improvement on other unseen languages.
-Table 3: Average number of different types of neurons for English and non-English languages of MistralMathOctopus on MGSM. We round the results to the nearest integer.
 Language Language-Specific Language-Related
@@
 non-English 613 2006
+Table 3: Average number of different types of neurons for English and non-English languages of MistralMathOctopus on MGSM. We round the results to the nearest integer.
+Further Analysis
+Uniqueness of English
+Since current LLMs are primarily pretrained on English data, English is often regarded as playing a special role within LLMs (Wendler et al.
+2024).
+In our work, we observe that English exhibits markedly different characteristics compared to other languages.
+Based on the identified neurons in our work, in Figure 2, deactivating the language neurons of English has a negligible impact on the model’s performance in English, which is entirely different from the behavior observed in other languages.
+This is consistent with the results of Tang et al.
+(2024).
+Furthermore, we quantify the sum of language-specific neurons and language-related neurons for English and non-English languages based on the MistralMathOctopus base model (Table 3).
+Our analysis reveals that English has significantly fewer neurons than other languages, both in terms of language-specific and language-related neurons.
+We hypothesize that this is due to that English actually possesses numerous language-related neurons.
+And since English serves as a pivot language, these language-related neurons are likely shared with almost all other languages.
+It causes them to be also regarded as general neurons.
+Variable (%) Language-Specific Language-Related
+Domain 80.7 92.3
+Alignment 95.6 92.1
 Table 4: Overlap ratio of different types of neurons across different domains and before and after alignment.
@@
 The results of MistralMathOctopus on MGSM are used as the fiducial value.
-Variable (%) Language-Specific Language-Related
-Domain 80.7 92.3
-Alignment 95.6 92.1
-4.7 Further Analysis
-Uniqueness of English
-Since current LLMs are primarily pretrained on English data, English is often regarded as playing a special role within LLMs (Wendler et al., 2024).
-In our experiments, we also observe that English exhibits markedly different characteristics compared to other non-English languages.
-Based on the identified neurons in our work, in Figure 2, we can find that deactivating the language neurons of English has a negligible impact on the model’s performance in English, which is entirely different from the behavior observed in other languages.
-This is also consistent with the results of Tang et al.
-(2024).
-Furthermore, based on this finding, we quantify the number of language neurons for English and non-English languages based on the MistralMathOctopus base model (Table 4.6).
-Our analysis reveals that English has significantly fewer neurons than other languages, both in terms of language-specific and language-related neurons.
-We hypothesize that this is due to the fact that English actually possesses numerous language-related neurons.
-And since English serves as a pivot language, these language-related neurons are likely shared with almost all other languages.
-It causes them to be confounded with language-agnostic neurons.
 Stability of Neuron Distributions
 We discuss the stability of neuron distributions across different data domains, as well as before and after alignment.
-To quantify the stability of neuron distributions, we compute the neuron overlap ratio in both settings, with the results summarized in Table 4.6.
-We can find that although the exact positions of a few language neurons may vary across different settings, the positional distribution of most language neurons remains stable.
+To quantify the stability of neuron distributions, we compute the neuron overlap ratio in both settings, with the results summarized in Table 4.
+We can find that although the exact positions of a few neurons may vary across different settings, the positional distribution of most neurons remains stable.
 This also indicates good reliability and generalization of the language neurons identified under fixed hyperparameters.
-5 Conclusion
+Conclusion
 In this work, we systematically investigate the multilingual alignment from the perspective of language neurons.
-We propose a new language neuron identification algorithm based on entropy and probability value, which detects the language-specific neurons, language-related neurons, and language-agnostic neurons in LLMs.
-The validity of the identified neurons is confirmed through deactivation ablation experiments.
+First, based on the defects of the existing binary classification methodology, we propose a ternary classification methodology, which defines the language-specific neurons, language-related neurons, and general neurons.
+Then we propose a corresponding language neuron identification algorithm, which detects the above different types of neurons in LLMs.
 Furthermore, we examine the multilingual alignment mechanism by analyzing the roles of different types of neurons.
@@
 Meanwhile, we find that alignment promotes a greater reliance on shared language-related neurons across languages, rather than on language-specific neurons.
-We also explore the phenomenon of spontaneous multilingual alignment.
-Additionally, we provide further analysis and more empirical results based on the preceding findings.
-Appendix A Limitations
-We provide insights and analysis results for multilingual alignment and multilingual capabilities of LLMs, which allows for a better understanding of multilingualism.
-Despite we have conducted a systematic investigation in our work, there are still some limitations waiting for research.
-Due to the limited resources, we only conduct experiments on two different models and two different datasets.
-We are willing to perform a more comprehensive analysis on different scenarios if more resources are available in the future.
-Additionally, we don’t perform a finer-grained analysis of neurons within the same layer in this work.
-We would like to explore this in our future work.
-Appendix B Implementation Details
+This also corresponds to the phenomenon of spontaneous multilingual alignment.
+Overall, we provide further analysis and valuable insights to better understand multilingual alignment and multilingual capabilities of LLMs.
+Acknowledgement
+We would like to thank the anonymous reviewers for their insightful comments.
+Shujian Huang is the corresponding author.
+This work is supported by National Science Foundation of China (No.
+62376116, 62176120), research project of Nanjing University-China Mobile Joint Institute (NJ20250038), the Fundamental Research Funds for the Central Universities (No.
+2024300507, 2025300390).
+Appendix A Implementation Details
 Our experiments are conducted on 4 NVIDIA RTX A6000 GPUs or 4 NVIDIA GeForce RTX 3090 GPUs.
-We use the TRL [von Werra et al., 2020] and DeepSpeed [Rasley et al., 2020] frameworks for preference alignment, and the vLLM engine Kwon et al.
-[2023] for inference.
-B.1 MAPO
+We use the TRL (von Werra et al.
+2020) and DeepSpeed (Rasley et al.
+2020) frameworks for preference alignment, and the vLLM engine (Kwon et al.
+2023) for inference.
+MAPO
 We employ the officially released scripts [^4] https://github.com/NJUNLP/MAPO to generate the preference data.
@@
 It takes 7.5 hours for one alignment.
-B.2 Language Neurons Identification
+Language Neurons Identification
 We spend 25 minutes calculating the activation probability of each neuron with MGSM dataset on a single NVIDIA GeForce RTX 3090 GPU.
@@
 After obtaining the activation probability, it takes 3 minutes to identify how many languages each selected neuron is related to.
-Appendix C Licenses for Used Assets
+Appendix B Licenses for Used Assets
 We list the names of the licenses for each asset we utilized in our work:
@@
   • vLLM: Apache-2.0
-Appendix D Deactivation Ablation Experiments
+Appendix C Deactivation Ablation Experiments
 For the MistralMathOctopus model, Figure 6 presents the results of the deactivation ablation experiments conducted on MGSM using the neuron identification algorithm proposed by Tang et al.
-[2024].
+(2024).
 Additionally, Figure 7 presents the results of the deactivation ablation experiments on the out-of-domain test set, MSVAMP.
 To verify the generalization of our identification algorithm, we also deploy the MetaMathOctopus to conduct the deactivation ablation experiment, with the result in Figure 8.
-Appendix E Layer-wise distribution of the different types of neurons
-In Figure 9(a) and 9(b), we separately display the layer-wise distribution of the different types of neurons of MistarlMathOctopus on MSVAMP and MetaMathOctopus on MGSM to validate the generalization of our finding discussed in §4.3.
-Appendix F Layer-wise Changes in the Number of Different Types of Neurons During Alignment
-[…]
-Appendix G Spontaneous Multilingual Alignment
+Appendix D Layer-wise distribution of the different types of neurons
+In Figure 9(a) and 9(b), we separately display the layer-wise distribution of the different types of neurons of MistarlMathOctopus on MSVAMP and MetaMathOctopus on MGSM to validate the generalization of our findings.
+Appendix E Layer-wise Changes in the Number of Different Types of Neurons During Alignment
+Figure 11 shows the same pattern as the results in the main text, thereby supporting its generalization.
+Furthermore, for each dataset, we examine the layer-wise changes in the number of different types of neurons in the process of alignment in Figure 1213 and 14.
+The notation $\text{ckpt-}x$” denotes the model checkpoint obtained after $x$ training steps.
+Appendix F Spontaneous Multilingual Alignment
 We report the accuracy of the base model on the MGSM and MSVAMP benchmarks in Table 5.
 The expected accuracy of random guessing is 50.0%.
-"X/Y $\Rightarrow$ T" indicates that languages X and Y are used for alignment.
+”X/Y $\Rightarrow$ T” indicates that languages X and Y are used for alignment.
 The best results for each language are highlighted.
 It is evident that models trained on multilingual translation data substantially outperform the original model across a wide range of languages, indicating that multilingual training significantly enhances the model’s multilingual capabilities.
+Appendix G Limitations
+We provide insights and analysis results for multilingual alignment and multilingual capabilities of LLMs, which allows for a better understanding of multilingualism.
+Despite we have conducted a systematic investigation in our work, there are still some limitations waiting for research.
+Due to the limited resources, we only conduct experiments on two different models and two different datasets.
+We are willing to perform a more comprehensive analysis on different scenarios if more resources are available in the future.
+Additionally, we don’t perform a finer-grained analysis of neurons within the same layer in this work.
+We would like to explore this in our future work.
 (a) Base - Language-specific Neurons
 (b) Aligned - Language-specific Neurons
-Figure 6: PPL changes of MistralMathOctopus on MGSM when deploying the algorithm proposed by Tang et al. [2024].
+Figure 6: PPL changes of MistralMathOctopus on MGSM when deploying the algorithm proposed by Tang et al. (2024).
 (a) Base - Language-specific Neurons
 (b) Aligned - Language-specific Neurons
-(c) Base - Language Neurons
-(d) Aligned - Language Neurons
+(c) Base - Language-specific & -related Neurons
+(d) Aligned - Language-specific & -related Neurons
 Figure 7: PPL changes of MistralMathOctopus on MSVAMP after deactivating language-specific neurons or language neurons.
@@
 “Aligned” indicates the results of the aligned model.
-(a) Base - Language Neurons on MGSM
-(b) Aligned - Language Neurons on MGSM
-Figure 8: PPL changes of MetaMathOctopus on MGSM after deactivating language neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model.
+(a) Base - MGSM
+(b) Aligned - MGSM
+Figure 8: PPL changes of MetaMathOctopus on MGSM after deactivating language-specific neurons and language-related neurons.
+“Base” indicates the results of the base model.
+“Aligned” indicates the results of the aligned model.
 (a) MistralMathOctopus on MSVAMP.
@@
 Figure 9: Layer-wise distribution of the different types of neurons.
+(a) Before alignment
+(b) After alignment
+Figure 10: Numbers of neurons shared by NN languages before and after alignment.
 Figure 11: Layer-wise changes in the number of MetaMathOctopus
@@
 (b) MSVAMP
-Figure 13: Layer-wise changes in the number of language-agnostic neurons of MistralMathOctopus during the alignment.
+Figure 14: Layer-wise changes in the number of general neurons of MistralMathOctopus during the alignment.
 Tested on MGSM bn th sw ja zh ru de es fr en Avg.